OpenStudy (anonymous):

Let x=log 2 and y=log 3 1) log 27

OpenStudy (anonymous):

It says express in terms of x and or y

OpenStudy (anonymous):

so your log is the decadic logarithm? the base is 10?

OpenStudy (anonymous):

I'm late, but I wanna step through this problem with properties of logs. \[ \begin{split} \log(27) &= \log(3\times 9) &(1)\\ &= \log(3\times 3 \times 3) &(2)\\ &= \log(3) + \log(3) + \log(3)& (3)\\ &= y + y + y & (4)\\ &=3y &(5) \end{split} \]The transition from step \((2)\) to step \((3)\) is just using the logarithmic identity: \[ \log(a\times b) = \log(a) + \log(b) \]

OpenStudy (anonymous):

Other logarithmic properties that is it good to know: \[ \log(a \times b) = \log (a) +\log(b) \\ \log(a \div b) = \log (a) -\log(b) \\ \log(a^b) = b\log (a) \\ \log(\sqrt[b]{a}) = \frac{1}{b}\log (a) \\ \log(1/a) = -\log (a) \\ \log_b(b) = 1\\ \log_b(1) = 0 \]